Demand for Orange Juice is given as
Qd = 5000 – 2500 P + 1200 I + 650 E – 255 Ps
Suppose Income is I = Rs.500, Expectations E = 55, and Price of Ps = Rs 25.
a. Find the Demand Equation.
b. Using the demand function from part a.,
Calculate Elasticity of Demand for price range of Rs.125 and Rs.155.
c. What will be the ‘Price Elasticity of Demand’ at P = Rs.125?
d. Interpret the Elasticity of Demand calculated in (C) above.
1
Expert's answer
2020-11-26T07:23:05-0500

a. If I = Rs.500, E = 55, Ps = Rs 25, then:

"Qd = 5000 \u2013 2500P + 1200\u00d7500 + 650\u00d755 \u2013 255\u00d725 = 634375 - 2500P."

b. Elasticity of Demand for price range of Rs.125 and Rs.155 is:

"Ed = \frac{321875 - 246875} {125 - 155} × \frac{125 + 155} {321875 - 246875} = - 9.33."

c. The ‘Price Elasticity of Demand’ at P = Rs.125 is:

"Ed = -2500\u00d7125\/321875 = 0.97."

d. The Demand calculated in (C) above is inelastic.